OpenStudy (anonymous):

solve for yfor both equations x-3y=3 & x+3y=9 just need to know if im right, compare answers?

OpenStudy (anonymous):

so for both of the equations get y by itself

OpenStudy (anonymous):

x-3y=3 -x -x 3y= -x +3 divide by -3 y= 1/3x ? thats my answer for the first equation x+3y=9 -x -x 3y= -x +9 divide by 3 y= -1/3+3 ? thats my answer for the second am i right? & how would i graph i this?

OpenStudy (anonymous):

got this for the first one y = (x-3)/3 and how did you get rid of the x on the second equation

OpenStudy (anonymous):



OpenStudy (anonymous):

the graph is for this equation :y= -1/3+3

OpenStudy (anonymous):

y = (9-x)/3 for second

OpenStudy (anonymous):

on the first equation you dropped the negative

OpenStudy (anonymous):

nvm you didnt my bad

OpenStudy (anonymous):

how do i graph the second equation?

OpenStudy (anonymous):

i did that :D

OpenStudy (anonymous):

i mean how do i graph both equations,?

OpenStudy (anonymous):

In the first equation: x - 3y = 3 Subtract x. -3y = -x + 3 Divide by -3. y = 1/3 - 1 To graph, remember that this is Slope-Intercept form. y = mx + b, where m is slope, and b is the y-intercept. First plot the y-intercept on the graph. y-intercept is where the line crosses the y-axis, so you would plot the point on (0, -1). Now is when the slope comes in. Remember slope is rise/run. rise = 1, run = 3. From (0, -1) RISE up 1 space, then RUN to the right 3 spaces. You'll end up on point (3, 0). Draw a line through points (0, -1) and (3, 0)

OpenStudy (anonymous):

one equation is y = 1/3x - 1 

OpenStudy (anonymous):

To graph #2, do the same. Plot the y-intercept and use slope(rise/run) to find another point. Then draw a line in between. y = -1/3 + 3

OpenStudy (anonymous):

For #2... The RISe will actually go down 1 space, since it is a -1.

OpenStudy (fools101):

Here try this Web site! www.mathway.com Let me know if this helps!